# On multi-graded Proj schemes

### v3

@@ 4.3 The Springer resolution
We continue with the setting of §4.2, in the special case when $\widetilde{V}$ is the Lie algebra $\mathbf{g}$ of $\mathbf{G}$ and $V$ is the Lie algebra $\mathbf{u}$ of the unipotent radical $\mathbf{U}$ of $\mathbf{B}$. In this case the vector bundle $\mathbf{G}\times^{\mathbf{B}}\mathbf{u}$ is called the Springer resolution, and usually denoted $\widetilde{\mathcal{N}}$. (The name is justified by the fact that, under appropriate technical assumptions, $\widetilde{\mathcal{N}}$ is a resolution of singularities of the nilpotent cone $\mathcal{N}\subset\mathbf{g}$ of $\mathbf{G}$, first introduced by Springer.) This varieties appears in numerous works in Geometric Representation Theory, among which [AB09] (in the case when $𝕜\Bbbk$ is an algebraic closure of the field $\mathbb{Q}_{\ell}$). In that reference the authors introduce an $\mathbb{X}$-graded ring, the spectrum of which is denoted $\hat{\tilde{\mathcal{N}}}_{af}$ in [AB09, §3.1]. This construction is reproduced in [AR, §6.2.2], where the spectrum is denoted $\widehat{\mathcal{N}}_{\mathcal{X}}$. The definition makes sense over an arbitrary algebraically closed field; here, we will denote this ring by $A_{\mathbf{u}}^{\prime}$. By construction, we have morphisms of $\mathbb{X}$-graded rings
@@ 5 Quasi-coherent sheaves on Proj schemes
In this section we study quasi-coherent shaves on Proj schemes, generalizing classical results for ${\operatorname{Proj}}$ schemes of ${\mathbb{N}}$-graded rings. Our presentation follows and extends [StP, Tag 01MJ] and [StP, Tag 01MM]. Other pioneering works on quasi-coherent sheaves on Brenner-Schröer Proj may be found in [MRo24].
@@ 5.1 Sheaves associated with graded modules
Let $Q$ be an $M$-graded $A$-module. For any homogeneous multiplicative subset $S\subset A$, we have considered in Definition 2.9 the $A_{(S)}$-module $Q_{(S)}$.
@@
There exists a unique quasi-coherent $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-module $\widetilde{Q}$ such that
$\Gamma\bigl{(}D_{\dagger}(S),\widetilde{Q}\bigr{)}=Q_{(S)}$
for every $S\in{\mathcal{F}}_{A}$.
Proof.
This is immediate by glueing. ∎
@@
An $M$-graded $A$-module $Q$ will be called negligible if $\tilde{Q}=0$. We will denote by ${\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}$ the full subcategory of ${\mathsf{Mod}}^{M}(A)$ whose objects are the negligible modules. Since the functor $Q\mapsto\widetilde{Q}$ is exact, this is a Serre subcategory, see [StP, Tag 02MQ], and our functor factors through an exact functor
@@
Proof.
The morphisms $Q_{0}\to Q_{(S)}$ given by $x\mapsto\frac{x}{1}$, where $S$ runs through ${\mathcal{F}}_{A}$, are compatible by restriction, hence glue to the desired morphism. ∎
@@
Let $P$, $Q$ be graded $A$-modules. There is a canonical map of $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-modules
@@
Constructing a morphism as displayed is equivalent to constructing an $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-bilinear map
$\widetilde{P}\times\widetilde{Q}\longrightarrow\widetilde{P\otimes_{A}Q},$
see [StP, Tag 01CA]. It suffices to define this map on sections over the opens $(D_{\dagger}(S):S\in{\mathcal{F}}_{A})$ compatible with restriction maps. On $D_{\dagger}(S)$, with $S=\{a_{i}:i\in I\}$, we use the $A_{(S)}$-bilinear map $P_{(S)}\times Q_{(S)}\to(P\otimes_{A}Q)_{(S)}$ given by $(x/a^{\nu},y/a^{\nu^{\prime}})\mapsto(x\otimes y)/a^{\nu+\nu^{\prime}}$. The commutation of the diagram follows from definitions. ∎
@@
Consider the setting of Remark 3.7, and denote by ${\mathsf{Mod}}^{M,H}(A)$ the category of $H$-equivariant $M$-graded $A$-modules. Let also ${\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}$ be the Serre subcategory consisting of objects which are negligible as $M$-graded $A$-modules. On the other hand, consider the category $\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$ of $H$-equivariant quasi-coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. (For a review of equivariant quasi-coherent sheaves, see e.g. [MRi16, Appendix A].) Then very similar considerations to those about allow to construct an exact functor
${\mathsf{Mod}}^{M,H}(A)\to\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$
which factors through the quotient ${\mathsf{Mod}}^{M,H}(A)/{\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}$.
@@ 5.2 Twisting sheaves
    The quasi-coherent sheaf $\widetilde{A(\alpha)}$ is called the $\alpha$-th-twist of the structure sheaf of $\mathrm{Proj}^{M}(A)$, and is denoted $\mathcal{O}_{\mathrm{Proj}^{M}(A)}(\alpha)$.
@@
a structure of $M$-graded sheaf of $\mathcal{O}_{\mathrm{Proj}^{M}(A)}$-algebras, and then an $M$-graded ring structure on
@@
    There is a canonical morphism of graded rings
@@ 5.3 Graded modules associated with sheaves
    The sheaf ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}$ is invertible, and in fact isomorphic to $({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})_{|D_{\dagger}(S)}$.
@@
Since $\alpha\in M[\underline{S}]$, there exists an invertible element $a\in A_{S}$ of degree $\alpha$. Then the map $x\mapsto a\cdot x$ induces an isomorphism of $A_{(S)}$-modules $A_{(S)}\cong(A(\alpha))_{(S)}$. We deduce (1).
Similarly, given a graded $A$-module $Q$, the map $q\mapsto a\cdot q$ induces an isomorphism of $A_{(S)}$-modules $Q_{(S)}\cong(Q(\alpha))_{(S)}$, which implies (2). The statement in (3) is the special case of (2) where $Q=A(\alpha^{\prime})$. Finally, (4) is a direct consequence of (3). ∎
@@
For any ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$, there exists a canonical (in particular, functorial) morphism
$\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$
in ${\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$.
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered (by definition) by the open subschemes $D_{\dagger}(S)$, to prove the lemma it suffices to construct morphisms
@@
for $S\in{\mathcal{F}}_{A}$, which coincide on intersections of such open subschemes. And given $S\in{\mathcal{F}}_{A}$, by basic properties of quasi-coherent sheaves on affine schemes (see [StP, Tag 01I7]), to define such a morphism it suffices to define a morphism of $A_{(S)}$-modules
@@
Now an element of $\Gamma_{\bullet}({\mathcal{Q}})_{(S)}$ can be represented by a fraction $\frac{m}{s}$ where $s\in S$ and $m$ is a homogeneous element of $\Gamma_{\bullet}({\mathcal{Q}})$ of degree $\deg(s)$, i.e. a global section of ${\mathcal{Q}}(\deg(s))$. Consider
@@
The element $\frac{1}{s}\in A_{S}$ has degree $-\deg(s)$, hence defines a section of ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(S)}(-\deg(s))$ on $D_{\dagger}(S)$, which we denote $s^{-1}$. Then the product $m_{|D_{\dagger}(S)}\otimes s^{-1}$ defines a section of
${\mathcal{Q}}(\deg(s))\otimes_{{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}}{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(-\deg(s))\cong{\mathcal{Q}}$
over $D_{\dagger}(S)$. (Here the isomorphism is provided by Proposition 5.8(4).)
@@ 5.4 Maximally relevant families
    Assume that $M=\mathbb{Z}$, and that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra. (In particular, we are in the $\mathbb{N}$-graded case.) If $(a_{i}:i\in I)$ is a family of generators of $A_{1}$ as an $A_{0}$-module, since any relevant element must have positive degree the condition of Corollary 3.10 is satisfied. Clearly, each $a_{i}$ is maximally relevant.
@@
(see (3.4)) is a Zariski trivial principal bundle for the group scheme $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$.
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered by the open subschemes $(D_{\dagger}(S):S\in{\mathcal{F}})$, we have $\widetilde{Q}=0$ iff $Q_{(S)}=0$ for any $S\in{\mathcal{F}}$. Now since ${\mathcal{F}}\subset{\mathcal{F}}^{\mathrm{max}}$, for any $S\in{\mathcal{F}}$ the ring $A_{S}$ has invertible elements of all degrees, so that $Q_{(S)}=0$ iff $Q_{S}=0$. Finally, it is clear from definitions that $Q_{S}=0$ iff for any $q\in Q$ there exists $s\in S$ such that $s\cdot q=0$. ∎
@@
To prove the proposition we need to define functorial morphisms
$\varepsilon_{{\mathcal{Q}}}:\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$
for ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$ and
$\eta_{Q}:Q\to\Gamma_{\bullet}(\widetilde{Q})$
for $Q\in{\mathsf{Mod}}^{M}(A)$, which satisfy the usual zigzag relations. Here $\varepsilon$ is provided by Lemma 5.9. (This does not require any assumption.) To define $\eta$, we observe that for $Q\in{\mathsf{Mod}}^{M}(A)$ we have
@@
where the isomorphism follows from Item (2) in Corollary 5.11. The desired morphism is therefore provided by Fact 5.7(2).
@@
Our assumptions imply that ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$ is a line bundle (see Corollary 5.11), so that ${\operatorname{Proj}}^{M}(A)_{\tilde{f}}$ is well defined. In view of our assumption, to prove the statement it suffices to prove that for any $S\in{\mathcal{F}}^{\mathrm{m}}_{A}$ we have
$D_{\dagger}(S)_{\tilde{f}_{S}}=D_{\dagger}(S)\cap D_{\dagger}(f),$
where $\tilde{f}_{S}$ is the restriction of $\tilde{f}$ to $D_{\dagger}(S)$. Fix such an $S$. As in the proof of Proposition 5.8(1), there exist $a,s\in S$ such that $\deg(a)-\deg(s)=\deg(f)$, and then multiplication by $\frac{a}{s}$ induces an isomorphism of sheaves
@@
The inverse image of $\tilde{f}_{S}$ under this isomorphism is $\frac{fs}{a}$, which implies that
$D_{\dagger}(S)_{\tilde{f}_{S}}={\operatorname{Spec}}\left((A_{(S)})_{\frac{fs}{a}}\right).$
@@
Assume that (5.8) is satisfied, and moreover that ${\operatorname{Proj}}^{M}(A)$ is quasi-compact. Then for any ${\mathcal{Q}}$ in $\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$, the morphism
$\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$
of Lemma 5.9 is an isomorphism.
@@
over the $\mathbb{N}$-graded ring
@@
and denoting by $\tilde{f}$ the image of $f$ in $\Gamma({\operatorname{Proj}}^{M}(A),{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$, then we have
$\Gamma_{\bullet}({\mathcal{Q}})_{(f)}=Q_{(\tilde{f})}.$
Since ${\operatorname{Proj}}^{M}(A)$ is quasi-compact (by assumption) and quasi-separated (see Lemma 3.3), by [StP, Tag 01PW] the right-hand side identifies with
$\Gamma({\operatorname{Proj}}^{M}(A)_{\tilde{f}},{\mathcal{Q}}),$
i.e. with $\Gamma(D_{\dagger}(s),{\mathcal{Q}})$ by Lemma 5.18, which finishes the proof. ∎
@@
Let $\mathsf{A}$, $\mathsf{B}$ be abelian categories, and let
$L:\mathsf{A}\to\mathsf{B},\quad R:\mathsf{B}\to\mathsf{A}$
be functors. Assume that
@@
Then $L$ factors through an equivalence of categories
$\overline{L}:\mathsf{A}/\ker(L)\xrightarrow{\sim}\mathsf{B},$
whose quasi-inverse is the composition of $R$ with the quotient functor $\mathsf{A}\to\mathsf{A}/\ker(L)$.

### v4

@@ 4.3 The Springer resolution
We continue with the setting of §4.2, in the special case when $\widetilde{V}$ is the Lie algebra $\mathbf{g}$ of $\mathbf{G}$ and $V$ is the Lie algebra $\mathbf{u}$ of the unipotent radical $\mathbf{U}$ of $\mathbf{B}$. In this case the vector bundle $\mathbf{G}\times^{\mathbf{B}}\mathbf{u}$ is called the Springer resolution, and usually denoted $\widetilde{\mathcal{N}}$. (The name is justified by the fact that, under appropriate technical assumptions, $\widetilde{\mathcal{N}}$ is a resolution of singularities of the nilpotent cone $\mathcal{N}\subset\mathbf{g}$ of $\mathbf{G}$, first introduced by Springer.) This variety appears in numerous works in Geometric Representation Theory, among which [AB09] (in the case when $𝕜\Bbbk$ is an algebraic closure of the field $\mathbb{Q}_{\ell}$). In that reference the authors introduce an $\mathbb{X}$-graded ring, the spectrum of which is denoted $\hat{\tilde{\mathcal{N}}}_{af}$ in [AB09, §3.1]. This construction is reproduced in [AR, §6.2.2], where the spectrum is denoted $\widehat{\mathcal{N}}_{\mathcal{X}}$. The definition makes sense over an arbitrary algebraically closed field; here, we will denote this ring by $A_{\mathbf{u}}^{\prime}$. By construction, we have morphisms of $\mathbb{X}$-graded rings
@@ 5 Quasi-coherent sheaves on Proj schemes
In this section we study quasi-coherent shaves on Proj schemes, generalizing classical results for ${\operatorname{Proj}}$ schemes of ${\mathbb{N}}$-graded rings. Our presentation follows and extends [StP, Tag 01MJ] and [StP, Tag 01MM]. Other pioneering works on quasi-coherent sheaves on Brenner–Schröer Proj schemes include [MRo24].
@@ 5.1 Sheaves associated with graded modules
Let $Q$ be an $M$-graded $A$-module. For any homogeneous multiplicative subset $S\subset A$, we have considered in Definition 2.9 the $A_{(S)}$-module $Q_{(S)}$. The following fact is immediate by glueing of quasi-coherent sheaves.
@@
There exists a unique quasi-coherent $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-module $\widetilde{Q}$ such that $\Gamma\bigl{(}D_{\dagger}(S),\widetilde{Q}\bigr{)}=Q_{(S)}$ for every $S\in{\mathcal{F}}_{A}$.
@@
An $M$-graded $A$-module $Q$ will be called negligible if $\widetilde{Q}=0$. We will denote by ${\mathsf{Mod}}^{M}(A)_{\mathrm{neg}}$ the full subcategory of ${\mathsf{Mod}}^{M}(A)$ whose objects are the negligible modules. Since the functor $Q\mapsto\widetilde{Q}$ is exact, this is a Serre subcategory, see [StP, Tag 02MQ], and our functor factors through an exact functor
@@
The following fact is again immediate, by glueing of sections of quasi-coherent sheaves.
@@
Let $P$, $Q$ be graded $A$-modules. There is a canonical morphism of quasi-coherent $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-modules
@@
Constructing a morphism as displayed is equivalent to constructing an $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}$-bilinear map $\widetilde{P}\times\widetilde{Q}\longrightarrow\widetilde{P\otimes_{A}Q}$, see [StP, Tag 01CA]. It suffices to define this map on sections over the opens $(D_{\dagger}(S):S\in{\mathcal{F}}_{A})$ compatible with restriction maps. On $D_{\dagger}(S)$, with $S=\{a_{i}:i\in I\}$, we use the $A_{(S)}$-bilinear map $P_{(S)}\times Q_{(S)}\to(P\otimes_{A}Q)_{(S)}$ given by $(\frac{x}{a^{\nu}},\frac{y}{a^{\nu^{\prime}}})\mapsto\frac{x\otimes y}{a^{\nu+\nu^{\prime}}}$. The commutation of the diagram follows from definitions. ∎
@@
Consider the setting of Remark 3.8, and denote by ${\mathsf{Mod}}^{M,H}(A)$ the category of $H$-equivariant $M$-graded $A$-modules. Let also ${\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}$ be the Serre subcategory consisting of objects which are negligible as $M$-graded $A$-modules. On the other hand, consider the category $\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$ of $H$-equivariant quasi-coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. (For a review of equivariant quasi-coherent sheaves, see e.g. [MRi16, Appendix A].) Then very similar considerations to those above allow to construct an exact functor ${\mathsf{Mod}}^{M,H}(A)\to\mathsf{QCoh}^{H}({\operatorname{Proj}}^{M}(A))$ which factors through the quotient ${\mathsf{Mod}}^{M,H}(A)/{\mathsf{Mod}}^{M,H}(A)_{\mathrm{neg}}$.
@@ 5.2 Twisting sheaves
    The quasi-coherent sheaf $\widetilde{A(\alpha)}$ on ${\operatorname{Proj}}^{M}(A)$ is denoted $\mathcal{O}_{{\operatorname{Proj}}^{M}(A)}(\alpha)$.
@@
a structure of $M$-graded sheaf of $\mathcal{O}_{\mathrm{Proj}^{M}(A)}$-algebras, and an $M$-graded ring structure on
@@
    There is a canonical morphism of $M$-graded rings
@@ 5.3 Graded modules associated with sheaves
    The sheaf ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\alpha)_{|D_{\dagger}(S)}$ is invertible, and in fact isomorphic to ${\mathcal{O}}_{D_{\dagger}(S)}$.
@@
Since $\alpha\in M[\underline{S}]$, there exists an invertible element $a\in A_{S}$ of degree $\alpha$. Then the map $x\mapsto a\cdot x$ induces an isomorphism of $A_{(S)}$-modules $A_{(S)}\cong(A(\alpha))_{(S)}=(A_{S})_{\alpha}$. We deduce (1). Similarly, given a graded $A$-module $Q$, the map $q\mapsto a\cdot q$ induces an isomorphism of $A_{(S)}$-modules $Q_{(S)}\cong(Q(\alpha))_{(S)}=(Q_{S})_{\alpha}$, which implies (2). The statement in (3) is the special case of (2) where $Q=A(\alpha^{\prime})$. Finally, (4) is a direct consequence of (3). ∎
@@
For any ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$, there exists a canonical (in particular, functorial) morphism $\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$ in ${\mathsf{Mod}}({\mathcal{O}}_{\mathrm{Proj}^{M}(A)})$.
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered (by definition) by the open subschemes $D_{\dagger}(S)$ for $S\in{\mathcal{F}}_{A}$ not containing 00, to prove the lemma it suffices to construct morphisms
@@
for such $S$, which coincide on intersections of such open subschemes. And given $S$, by basic properties of quasi-coherent sheaves on affine schemes (see [StP, Tag 01I7]), to define such a morphism it suffices to define a morphism of $A_{(S)}$-modules
@@
Now an element of $\Gamma_{\bullet}({\mathcal{Q}})_{(S)}$ can be represented by a fraction $\frac{m}{s}$ where $s\in S$ and $\in\Gamma_{\bullet}({\mathcal{Q}})_{\deg(s)}$, i.e. $m$ is a global section of ${\mathcal{Q}}(\deg(s))$. Consider
@@
The element $\frac{1}{s}\in A_{S}$ has degree $-\deg(s)$, hence defines a section of ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(S)}(-\deg(s))$ on $D_{\dagger}(S)$, which we denote $s^{-1}$. Then the product $m_{|D_{\dagger}(S)}\otimes s^{-1}$ defines a global section of
${\mathcal{Q}}(\deg(s))_{|D_{\dagger}(S)}\otimes_{D_{\dagger}(S)}{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(-\deg(s))_{|D_{\dagger}(S)}\cong{\mathcal{Q}}_{|D_{\dagger}(S)},$
wjere the isomorphism is provided by Proposition 5.8(4).
@@ 5.4 Maximally relevant families
The same comments as for (3.3) show that this condition is equivalent to the property that ${\operatorname{Proj}}^{M}(A)$ is covered by the open subschemes $D_{\dagger}(f)$ with $f$ maximally relevant.
@@
    Assume that $M=\mathbb{Z}$, and that $A$ is generated by $A_{1}$ as an $A_{0}$-algebra. (In particular, we are in the $\mathbb{N}$-graded case.) If $(a_{i}:i\in I)$ is a family of generators of $A_{1}$ as an $A_{0}$-module, since any non nilpotent relevant element must have positive degree the condition of Corollary 3.10 is satisfied. Clearly, each $a_{i}$ is maximally relevant.
@@
(see (3.4)) is a Zariski locally trivial principal bundle for the group scheme $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$.
@@
Since ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $(D_{\dagger}(S):S\in{\mathcal{F}})$, we have $\widetilde{Q}=0$ iff $Q_{(S)}=0$ for any $S\in{\mathcal{F}}$. Now since ${\mathcal{F}}\subset{\mathcal{F}}^{\mathrm{max}}$, for any $S\in{\mathcal{F}}$ the ring $A_{S}$ has invertible elements of all degrees, so that $Q_{(S)}=0$ iff $Q_{S}=0$. Finally, it is clear from definitions that $Q_{S}=0$ iff for any $q\in Q$ there exists $s\in S$ such that $s\cdot q=0$. ∎
@@
To prove the proposition we need to define functorial morphisms $\varepsilon_{{\mathcal{Q}}}:\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$ for ${\mathcal{Q}}\in{\mathsf{Mod}}({\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)})$ and $\eta_{Q}:Q\to\Gamma_{\bullet}(\widetilde{Q})$ for $Q\in{\mathsf{Mod}}^{M}(A)$, which satisfy the usual zigzag relations. Here $\varepsilon$ is provided by Lemma 5.9. (This does not require any assumption.) To define $\eta$, we observe that for $Q\in{\mathsf{Mod}}^{M}(A)$ we have
@@
where the isomorphism follows from Corollary 5.11(2). The desired morphism is therefore provided by Fact 5.7(2).
@@
We can assume that $f$ is nonzero. Our assumptions imply that ${\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$ is a line bundle (see Corollary 5.11), so that ${\operatorname{Proj}}^{M}(A)_{\tilde{f}}$ is well defined. In view of our assumption, to prove the statement it suffices to prove that for any $S\in{\mathcal{F}}^{\mathrm{m}}_{A}$ we have
$D_{\dagger}(S)_{\tilde{f}_{|D_{\dagger}(S)}}=D_{\dagger}(S)\cap D_{\dagger}(f).$
Fix such an $S$. As in the proof of Proposition 5.8(1), there exist $a,s\in S$ such that $\deg(a)-\deg(s)=\deg(f)$, and then multiplication by $\frac{a}{s}$ induces an isomorphism of sheaves
@@
The inverse image of $\tilde{f}_{|D_{\dagger}(S)}$ under this isomorphism is $\frac{fs}{a}$, which implies that
$D_{\dagger}(S)_{\tilde{f}_{|D_{\dagger}(S)}}={\operatorname{Spec}}\left((A_{(S)})_{\frac{fs}{a}}\right).$
@@
Assume that (5.8) is satisfied, and moreover that ${\operatorname{Proj}}^{M}(A)$ is quasi-compact. Then for any ${\mathcal{Q}}$ in $\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$, the morphism $\widetilde{\Gamma_{\bullet}({\mathcal{Q}})}\to{\mathcal{Q}}$ of Lemma 5.9 is an isomorphism.
@@
over the $\mathbb{Z}$-graded ring
@@
and denoting by $\tilde{f}$ the image of $f$ in $\Gamma({\operatorname{Proj}}^{M}(A),{\mathcal{O}}_{{\operatorname{Proj}}^{M}(A)}(\deg(f))$, then we have $\Gamma_{\bullet}({\mathcal{Q}})_{(f)}=Q_{(\tilde{f})}$. Since ${\operatorname{Proj}}^{M}(A)$ is quasi-compact (by assumption) and quasi-separated (see Lemma 3.4), by [StP, Tag 01PW] the right-hand side identifies with $\Gamma({\operatorname{Proj}}^{M}(A)_{\tilde{f}},{\mathcal{Q}})$, i.e. with $\Gamma(D_{\dagger}(f),{\mathcal{Q}})$ by Lemma 5.18, which finishes the proof. ∎
@@
Let $\mathsf{A}$, $\mathsf{B}$ be abelian categories, and let $L:\mathsf{A}\to\mathsf{B}$ and $R:\mathsf{B}\to\mathsf{A}$ be functors. Assume that
@@
Then $L$ factors through an equivalence of categories $\overline{L}:\mathsf{A}/\ker(L)\xrightarrow{\sim}\mathsf{B}$, whose quasi-inverse is the composition of $R$ with the quotient functor $\mathsf{A}\to\mathsf{A}/\ker(L)$.